### v1

# Unsupervised deep learning model for fast energy layer pre-selection of delivery-efficient proton arc therapy plan optimization of nasopharyngeal carcinoma

Abstract
Objective. Proton arc therapy (PAT) is an emerging and promising modality in radiotherapy, offering several advantages over conventional intensity-modulated proton therapy (IMPT). However, identifying the optimal energy layer (EL) sequence remains computationally intensive due to the large number of possible energy layer transitions. This study proposes an unsupervised deep learning framework for fast and effective EL pre-selection, aiming to minimize energy layer switch time while preserving high plan quality. Approach. We introduce a novel data representation method—spot-count representation—which encodes the number of proton spots intersecting the target and organs at risk (OARs) in a matrix structured by sorted gantry angles and energy layers. This representation is the input of a U-Net-based architecture, SPArcdl, which is trained to optimize a tri-objective function: maximizing target coverage, minimizing OAR exposure, and reducing energy switching time. The model is evaluated on $54$ nasopharyngeal cancer cases, and its performance is benchmarked against plans generated by SPArcparticle swarm. Main results. SPArcdl produces EL pre-selection that significantly improves both plan quality and delivery efficiency. Compared to SPArcparticle swarm, it enhances the conformity index by $0.16$ ($p<0.01$), reduces the homogeneity index by $0.71$ ($p<0.01$), shortens the energy switching time by $38.4\%$ ($p<0.01$), and lowers the mean dose to brainstem by $0.21$ ($p<0.01$). The results unintentionally reveal employing unchanged ELS is more time-wise efficient than descended ELS. SPArcdl’s inference time is within 1 second. Significance. SPArcdl is a fast and effective tool for generating high-quality PAT plans by strategically pre-selecting energy layers to reduce delivery time while maintaining excellent dosimetric performance. The proposed spot-count representation lays a foundation for incorporating unsupervised deep learning approaches into the optimization of proton arc therapy planning.
  • Keywords: Proton arc therapy; spot scanning; beam delivery time; energy layer selection; deep learning
1 Introduction
Proton arc therapy (PAT) is an advanced and promising radiation therapy modality that combines the physical advantages of proton therapy with continuous rotational beam delivery. Since its initial introduction by Sandison et al. [20] in 1997, PAT has seen renewed interest and technological advancement with the emergence of pencil beam scanning (PBS) and the development of Spot-Scanning Proton Arc Therapy (SPArc) by Ding et al. [6] in 2016. Several studies have suggested that, in comparison to traditional methods like intensity-modulated proton therapy (IMPT), PAT may provide improved dose distribution [13], greater treatment robustness [11, 4], enhanced radio-biological effectiveness [1, 3], and a more streamlined treatment process [12]. Clinical and simulation studies have since demonstrated PAT’s potential benefits across various tumor sites, including prostate, lung, breast, brain, and head and neck cancers [5, 7, 11, 4, 10, 21, 19].
A major approach to solve the PAT problem is to break it down into two steps: pre-selecting ELs in advance, then using the selected ELs to complete regular spot weight optimization [27]. A significant challenge to the clinical adoption of PAT lies in the optimization of energy layer (EL) sequences. In SPArc, the system is required to deliver protons at selected energy layers on each gantry angle. With hundreds of control points involved, frequent switching between energy layers can lead to substantial delays in treatment delivery. This is especially true for upward energy switches, which are time-intensive due to the magnetic hysteresis in the proton therapy system [8, 30, 2, 23, 22]. Consequently, minimizing energy layer switching (ELS) time without compromising dosimetric quality has become a focus in PAT plan optimization.
Several EL sequence optimization approaches have been proposed to address this challenge. Sanchez-Parcerisa et al. [19] proposes and compares four range optimization algorithms to select optimal mono- and bi-energetic beam energies for proton modulated arc therapy (PMAT), aiming to minimize energy-switching time while maintaining dosimetric quality. However, this method shows constraints in performance when applied to intricate or irregular target geometries. The SPArc algorithm [6] improves dose delivery by iteratively refining gantry angles and redistributing ELs. As a subsequent method, SPArcseq [12] reduces switching time by aligning energy sequences with the gantry’s rotational direction, and SPArcparticle swarm (SPArcps) [17] builds upon it by segmenting the arc into sectors based on mean water equivalent thickness (WET) and selecting ELs that minimize switching time across control points. However, SPArcps may need more objective functions on organs at risk (OARs) to optimize Head and Neck cancer, including nasopharyngeal carcinoma (NPC). Additionally, the ELO-SPAT framework [28] integrates multiple objectives, including dose fidelity, energy sequencing constraints, and regularization terms, into a unified mathematical formulation that penalizes upward energy switches while exploiting downward transitions to enhance delivery efficiency. However, the optimization time reported in these studies [29, 28, 15]—ranging from several minutes to multiple hours—could pose a barrier to the clinical adoption of the framework. Based on the ground truth provided by modified ELO-SPAT, Ma et al. [15] demonstrates the feasibility of using supervised machine learning method to integrate features in geometric, WET and beamlet and generate ELs for each field. However, a supervised approach not only adds to the development workload due to ground truth generation but also constrains model performance to the quality of the pre-existing algorithms.
A gap in the literature is the lack of deep learning-based approaches that can swiftly and directly predict optimal EL sequences, particularly using unsupervised learning. Deep learning has demonstrated strong potential in radiation oncology, offering high efficiency, automation, and generalization capacity [9]. Yet, its application to PAT planning, and especially to EL pre-selection, remains largely unexplored. Unsupervised learning offers a key advantage in this context: it eliminates the dependency on manually labeled data or algorithm-generated ground truths, enabling researchers to directly embed domain knowledge into customized loss functions that reflect clinical priorities such as dose coverage, OAR sparing, and delivery efficiency.
In this study, we propose SPArcdl, an unsupervised deep learning model for fast and effective energy layer pre-selection in PAT plan optimization, which generates a data format specifically designed to support the application of deep learning models in EL pre-selection. These matrices serve as proxies for dose coverage and are analyzed by a U-Net-based neural network. The model is trained end-to-end using a tri-objective loss function that (1) maximizes target coverage, (2) minimizes OAR exposure, and (3) penalizes frequent and time-consuming EL switches.
We evaluate SPArcdl on 54 NPC cases and compare its performance with SPArcps. Results show that SPArcdl significantly EL switch time—by 38.4%—while enhancing conformity index by $0.16$ ($p<0.01$), lowering homogeneity by $0.71$ ($p<0.01$), reducing the mean dose of brainstem by $0.21$ ($p<0.01$). By bypassing the need for supervised labels and offering rapid inference within $1$ second, SPArcdl demonstrates a viable path toward integrating deep learning into real-time PAT planning workflows, supporting broader clinical adoption through enhanced efficiency, automation, and plan quality.
2 Materials and methods
2.1 Dataset and spot-count data representation
The treatment planning CT scans, acquired with a Philips Big Bore CT, of $54$ patients who has undergone nasopharyngeal carcinoma (NPC) radiotherapy were collected from our collaborative hospital [^1] Union Hospital, Tongji Medical College, Huazhong University of Science and Technology. Our method for pre-selecting ELs depends on the open-source software matRad (Department of Medical Physics in Radiation Oncology, German Cancer Research Center-DKFZ) [25] which uses the CTs to generate the initial complete range of ELs that longitudinally encompassed the entire treatment target at all gantry angles.
The inclusion of diverse information—such as beamlet, WET and geometric-related features [15]—may complicate the input structure of deep learning model and lead researchers to ambiguity. This study presents a novel data restructuring approach, termed spot-count representation, which constructs a data format specifically tailored to facilitate the application of deep learning models in EL pre-selection. The spot-count representation is founded on the assumption that the number of ELs, denoted by $E$, and the number of gantry angles, denoted by $G$, are finite. Accordingly, we construct a matrix $M\in\mathbb{R}^{G\times E}$, where rows correspond to sorted gantry angles and columns to sorted ELs, to record the number of proton spots intersecting a given anatomical structure. In this context, each element of $M$ represents the count of spots that intersect a specific structure $S$, organized by gantry angle and EL. For example, the number of spots intersecting CTV can be integrated into $M_{CTV}$.
Building on the concept of $M$, the number of spots intersecting one anatomical structure ($S_{1}$) but not another ($S_{2}$) can be represented by the element-wise subtraction $M_{S_{1}}-M_{S_{2}}$, resulting in a new matrix denoted as $M_{S_{1}-S_{2}}$. For instance, the number of spots that pass through the CTV while avoiding the brainstem is captured by the matrix $M_{CTV-Brainstem}$ ($M_{CTV-BRS}$). Since $M$ systematically records the number of spot intersections with anatomical structures, this data restructuring method is referred to as the spot-count representation.
Figure 1: (a) SPArcdl model architecture. (b) ELS time table displays the time cost when the gantry angle transitions from $g-1$ to $g$.
2.2 Method
We approximate the dose coverage of a volume as the number of spots intersection with the volume. Based on this approximation, matrix $M_{CTV}$ presents the dose coverage of CTV and $M_{CTV-BRS}$ presents dose covering CTV while dodging brainstem. Both matrices are formed and structured by gantry angles and EL. Previous method has found relative mean WET is key to mitigate the need for EL ascending and therefore improve delivery efficiency [17]. As shown in Figure 1, the mean value of WET for each gantry angle is calculated and fused with $M_{CTV}$ and $M_{CTV-BRS}$ by the gantry angle’s channel. These two matrices are then concatenated into a 2-channels tensor and fed to the network, $h$. In our experiments, the input is of size $2\times 72\times 103$, but the network is not limited to this specific size. This input size happens to be similar to the dimensionality of an image, which is channel $\times$ height $\times$ width. In the past decade, many deep learning architecture have been developed to solve computer vision tasks. We use $h$, a network similar to the classic image segmentation method U-Net  [18], to analyze the spot counts of CTV and brainstem in the context of ELs and gantry angles. The model consists of encoder and decoder sections with skip connection. We use convolution layers with a kernel size of $3$ and a stride of $1$ in both the encoder and decoder stages. Each convolutional layer is followed by a LeakyReLU activation function with a negative slope of $0.2$. The convolutional layers capture hierarchical spot-count distribution features of CTV and brainstem to estimate the optimal ELs. The encoder employs average pooling to reduce the spatial dimensions in half at each level. During the decoding stage, we alternate between upsampling, applying convolutions, and concatenating skip connections that transfer features learned in the encoding stages directly to the layers responsible for generating the EL selection. The spatial resolution of the encoder levels is one half of the previous level, and that of the decoder levels is twice of the previous level. Four levels in total are included in the network with depth of $16$, $32$, $64$, and $128$. We set $1$ as the channel number of network’s output and then perform Softmax on the logit vectors in gantry angle’s direction to obtain a probability tensor $P\in\mathbb{R}^{G\times EL}$. At inference stage, the EL corresponding to the highest probability in the logit vector for each gantry angle is selected as the final EL for that angle. We name this model, including the loss function, SPArcdl.
The dataset is divided randomly into $5$ folds for cross-validation, and the reported results represent the mean validation performance across all folds. The proposed model is trained end-to-end, optimized with the AdamW [14] optimizer using a learning rate of $10^{-5}$ and a batch size of $24$. Training is conducted for $1000$ epochs. The model is implemented in PyTorch[16] and trained on a single NVIDIA GeForce 3090 GPU, and inference is run on a 11th Gen Intel(R) i7 CPU.
2.3 Unsupervised Loss Function
The three objectives of SPArcdl are to (1) maximize the dose coverage on CTV; (2) minimize the dose coverage on brainstem; and (3) reduce the energy shifting time across all gantry angles. The EL pre-selection task can be formulated as an EL classification problem at all gantry angles. As a popular loss function of classification problems, cross entropy provides a well-founded, interpretable, and computationally efficient way to measures the difference between two probability distributions. In this paper, the in-used cross entropy loss is formulated in Eq.(1).
$CE(M,P)=-\frac{1}{NG}\Sigma_{n=1}^{N}\Sigma_{g=1}^{G}\Sigma_{el=1}^{EL}(M_{g,el}\cdot log(P_{g,el})),$ (1)
where $CE$ denotes cross entropy, $N$ denotes the number of patients, $EL$ denotes the number of energy layers, $G$ denotes the number of gantry angles, $M_{g,el}$ denotes the true labels for energy $el$ for angle $g$, and $P_{g,el}$ is the predicted probability for energy $el$ for angle $g$.
Specifically, to achieve objective (1), the first part of the loss function of Eq.(4), i.e. $CE(b(M_{CTV}),P)$, aims to shorten the distance between probability distribution of $P_{g}$ and $b(M_{CTV_{g}})$, so as to select the EL at angle $g$ that maximize the number of spots, or dose coverage, $g\in G$. $b$ is binarized-max function, which set the maximum value, or the maximum spot count, of each gantry angle as $1$ and other values as 00. Like-wise, to achieve objective (2), the second part of Eq.(4), i.e. $CE(b(M_{CTV-BRS}),P)$, is designed to maintain the coverage of CTV while reducing the brainstem dose exposure. The third part of Eq.(4) is introduced to penalize the ELS time, which can be generated using the ELS time table ($TB_{els}$) and formularized as Eq.(2).
$L_{time}=\frac{1}{NG}\Sigma_{n=1}^{N}\Sigma_{g=1}^{G}T(P_{g-1},P_{g},TB_{els}),$ (2)
where $T(P_{g-1},P_{g},TB_{els})$ is the extraction of ELS time when the gantry angle switches from $g-1$ to $g$ given the time table $TB_{els}$. Figure 1 (b) illustrates $TB_{els}$ where both the vertical (top to bottom) and horizontal (left to right) axes represent EL indices sorted in ascending order based on EL values. Both dimensions have a length of 102, reflecting the total of $102$ distinct ELs presented in the dataset used. The cell value of $TB_{els}$ represents the ELS time between two adjacent gantry angles, $g_{-1}$ and $g$. It can be observed from Figure 1 (b) the time cost of ascended ELS is higher than that of descended ELS. Note the ELS time is zero when EL is unchanged. The total loss function is
$L_{total}$ $=\gamma\cdot CE_{CTV}+\zeta\cdot CE_{CTV-BRS}+\upsilon\cdot L_{time}$ (3)
$=\gamma\cdot CE(b(M_{CTV}),P)+\zeta\cdot CE(b(M_{CTV-BRS}),P)+\upsilon\cdot L_{time},$ (4)
where $BRS$ denotes brainstem, $\gamma,\zeta$ and $\upsilon$ are hyper-parameters that can be tuned to ensure all components of the loss function are balanced. Through extensive experiments, we tune $\gamma$, $\zeta$ and $\upsilon$ to be 0.5, 0.4 and 0.1, respectively.
2.4 Evaluation
A total of $54$ patients who underwent radiotherapy for NPC were retrospectively selected for evaluation by SPArcps and SPArcdl. The performance of SPArcdl across all five validation datasets ($54$ cases) were aggregated for assessment. We calculate metrics mean for quantitative evaluation and select four representative cases for qualitative evaluation. Two structures, CTV and brainstem, are selected for evaluation. Full arc gantry angle selection was employed, with an angular sampling interval of $5$ degrees due to the constraints of computing power. Our energy selection method depended on matRad supplying the initial complete set of ELs that span the entire treatment target longitudinally for every gantry angle. A spot spacing of $5$ mm was used in both the beam and lateral directions. Dose calculations were performed with a resolution of $3\times 3\times 3$ $mm^{3}$, and the voxel size of the CT images was $5$ $cm^{3}$. The prescription dose of CTV is $54$ Gy with $30$ fractions, making $1.8$ Gy per fraction.
2.4.1 Plan quality evaluation
The performance of SPArcdl in respect to plan quality of the PAT plan generated by the predicted ELs is evaluated. The mean value of D98, D2, homogeneity index (HI), conformity index (CI), mean and max dose of brainstem are calculated for comparison between SPArcdl and SPArcps.
HI quantifies how uniformly the prescribed radiation dose is distributed within the target volume. A low HI value, ideally close to 00, indicating a more homogeneous dose distribution, suggesting less dose variation within the target. Eq.(5) presents the formula of HI described in [26].
$HI=\frac{D_{2}-D_{98}}{D_{prescription}},$ (5)
where D2 and D98 denotes the minimum dose received by 2% and 98% of the CTV, and Dprescription refers to the prescription dose. CI measures how well the high-dose region conforms to the shape and size of the target volume. A CI value close to $1$ indicates a well-conformed treatment plan. Eq.(6) presents the formula of CI described in [24].
$CI=\frac{TV_{95}}{TV}\times\frac{TV_{95}}{V_{95}},$ (6)
where TV95 denotes the portion of the target volume receiving $95\%$ of the prescription dose, V95 represents the total volume receiving at least $95\%$ of the prescription dose, and TV refers to the entire target volume. The mean and max dose of brainstem are reported to evaluate the OAR sparing of the proposed method. A lower amount of dose to brainstem represents a better brainstem protection in PAT. The qualitative evaluation to visualize plan quality is presented by the CT scans with dose distribution.
A plan of a previous machine learning approach for EL prediction [15] reported a limited dose conformity index with regard to relatively small tumor sizes. Therefore, the plan quality metrics generated by the proposed deep learning approach is also assessed in respect to gross tumor volume (GTV).
Figure 2: Left column are EL prediction of $4$ representative cases generated by SPArcdl. The horizontal axes represents gantry angles and vertical axes represents proton energy (Mev). Right column are CT slices with dose distribution of the corresponding cases.
2.4.2 EL switch time and model efficiency analysis
A commonly emphasized objective in existing studies is the reduction of treatment delivery time [27, 17, 12, 15, 29]. The duration of arc delivery is influenced by various machine-specific factors, such as the number and order of ELs, spot positions, and gantry angles. Among these, ELS time often contributes the most to overall delivery time, especially when the number of energy layers and energy switch-ups is not minimized [27, 17, 12]. The ELS time is evaluated and analyzed for both SPArcps and SPArcdl. The efficiency to generate EL pre-selection, or the inference time of SPArcdl, is also reported.
3 Result
3.1 Prediction
Figure 2 demonstrates the ELs of four representative cases generated by SPArcdl and the CT slices with dose distribution of the corresponding cases. While most of the predicted ELs are in the range between $80$ and $145$, many short horizontal lines can be observed, representing the ELs at these consecutive gantry angles are unchanged. The CT slices with dose distribution demonstrate a decent CTV coverage while maintaining acceptable sparing of brainstem.
Table 1: The dosimetric performance comparison of 54 cases of SPArcps and SPArcdl; D98: the minimum doses received by $98\%$ of the CTV; D2: the minimum doses received by $2\%$ of the CTV; CI: CTV conformity; HI: CTV homogeneity; Mean and Max: the mean dose and max dose for brainstem.
Structure Type SPArcps SPArcdl (Ours)
CTV D98 1.10$\pm$0.59 1.66$\pm$0.03
D2 1.98$\pm$0.13 1.83$\pm$0.01
CI 0.60$\pm$0.12 0.76$\pm$0.04
HI $\downarrow$ 0.88$\pm$0.70 0.17$\pm$0.04
BrainStem Mean $\downarrow$ 0.84$\pm$0.24 0.63$\pm$0.19
Max $\downarrow$ 1.94$\pm$0.16 1.79$\pm$0.11
3.2 Plan quality
Table 1 demonstrates the quantitative dosimetric performance comparison, including D98, D2, CI, HI, brainstem mean and max, of SPArcdl and SPArcps on nominal distribution. With regard to CTV, SPArcdl plan presents a significantly ($p<0.01$) smaller average HI compared to the SPArcps with average differences of $0.71$. The D2 values are close between the two plans with SPArcdl plan displaying a lower value, and this difference is statistically significant ($p<0.01$). These results indicate SPArcdl can provide a more homogeneous dose distribution. Furthermore, SPArcdl exhibits a higher CI value, differing by $0.16$ compared to SPArcps. The CI difference is significant ($p<0.01$), reflecting better conformity of SPArcdl to the target. In respect to brainstem, SPArcdl yields a lower mean and max dose, differing by $0.21$ and $0.15$. The mean and max difference are both statistically significant ($p<0.01$), demonstrating better sparing of SPArcdl on this critical structure.
A more intuitive performance comparison of SPArcdl (blue triangle) and SPArcps (red squares) plans is demonstrated in Figure 3, in which the distribution of CI, HI, D98, D2, mean and max dose of brainstem are presented as a function of GTV using kernel density estimates and scatter plots. In plot (a), the HI values of SPArcps plan range from about $0.2$ to $2.5$, with a dense cluster around 0.2-1, while those of SPArcdl plan are lower and with a tighter cluster, mostly between $0.1$ and $0.3$ suggesting improved homogeneity of the proposed method. Also in plot (c), the D98 values of SPArcdl plan are higher, mostly between $1.5$ and $1.8$, with a dense cluster around 1.6-1.7, indicating better minimum dose coverage. The CI values of SPArcdl are higher in plot (b), mostly between $0.6$ and $0.8$, with a dense cluster around the same range, showing better conformity to the target volume. The observations and findings of Figure 3 align with the insights of Table 1. In all four plots, the relatively denser cluster generated by SPArcdl plan indicates greater consistency of its performance.
The denser clustering and reduced variance observed in SPArcdl’s results in Figure 3 demonstrate its superior capacity to generate consistent outcomes compared to SPArcps in EL pre-selection.
Figure 3: Comparison of various metrics performance of $54$ cases between SPArcdl and SPArcps plans. Each diagram shows the relationship between GTV volume and the respective metric, displaying kernel density estimate (KDE) plots and scatter points for both methods.
A comparable plan in a previous study [15] reported low dose conformity index with respect to small lung tumor sizes. Although the primary focus of this study is NPC, we are interested in checking if we have the similar finding. As demonstrated in Figure 3 (b), the CI exhibits modest decline as the GTV volume reduces from approximately $100$ to $20$ cc. In short, SPArcdl demonstrates a slightly lower CI for relatively small NPC GTV volume.
3.3 EL switch time
The performance of the model is evaluated in terms of mean energy layer switch (ELS) time in Table 2. unchanged ELS refers to the scenario in which the EL remains unchanged when the beam angle transitions to the next position. In such cases, the ELS time is effectively zero, which is shorter than that required for both ELS upward or downward transitions. Table 2 presents the mean ELS associated with ascended, descended and unchanged transition of SPArcdl and SPArcps. The mean unchanged ELS of SPArcdl is $52.6$, which is dramatically higher than that of SPArcps or $0.3$. Compared to SPArcps, the number of descended ELS of SPArcdl is almost entirely shifted to the unchanged ELS. While the mean ascended ELS of SPArcdl is slightly higher than that of SPArcps, the notable advantage of SPArcdl lies in its higher frequency of unchanged ELS. This advantage not only offsets the the cost of slightly higher EL in ascended transitions but also contributes to energy switch time of $51.9$ seconds, an $38.4\%$ reduction from SPArcps. In addition, Figure 2 demonstrates a few cases of EL generated by SPArcdl, where many unchanged ELS can be observed.
Table 2: The mean energy transitions and ELS time of 54 cases.
SPArcps SPArcdl(Ours) Diff
Number ascended ELS 7.2 7.8 +0.6 (8.9%)
Number descended ELS 63.5 10.6 -52.9 (-83.4%)
Number unchanged ELS 0.3 52.6 +52.3 (158.5 times)
ELS time (sec) $\downarrow$ 84.3 51.9 -32.4 (-38.4%)
3.4 Model Efficiency
The mean inference time of SPArcdl for generating ELs pre-selection during SPArc plan optimization is $0.05$ seconds on a CPU, representing a significant improvement in speed compared to the performance of the previous method [17], which is $16$ seconds ($p<0.01$). This enhanced efficiency is primarily due to SPArcdl’s capability to simultaneously model the spot-organ and spot-target information across all gantry angles and produce EL pre-selection in a single forward pass.
4 Discussion
In this study, we proposed SPArcdl, a deep learning method to rapidly pre-select EL that maintain high treatment quality and significantly reduce ELS time in PAT, without relying on supervised training data. The proposed method is tested on $54$ NPC patient cases and evaluated against a published method SPArcps.
The data preprocessing involves matrices creation of $M_{CTV}$ and $M_{CTV-BRS}$, in which $M_{CTV}$ represents the number of proton spot intersecting the CTV and $M_{CTV-BRS}$ indicates the spots intersecting CTV but not the brainstem. The values of both both matrices are structured by gantry angles and EL. SPArcdl takes in these matrices and WET vectors to generate the optimal EL of all gantry angles. A tri-objective loss function is designed to encourage the number of spots on CTV, penalize the spots on brainstem and discourage time-consuming ELS. The plan quality of SPArcdl is comparable or better than that of SPArcps, and ELS time has been largely improved.
The contributions of this study is as follows. (1) To the best of our knowledge, the proposed SPArcdl is the first deep learning model solving EL pre-selection problem for PAT plan optimization. Compared to SPArcps, SPArcdl demonstrates comparable or superior dosimetric performance, including dose conformity, homogeneity, and OAR sparing, while achieving greater consistency in results. (2) The experimental results of SPArcdl unintentionally revealed that deploying high frequency of unchanged ELS can be more time-wise efficient than high frequency of descended ELS. SPArcdl achieves improvements in delivery efficiency by leveraging unchanged ELS to significantly reduce ELS time. (3) This study introduces a novel data representation, termed spot-count representation, which is used as a proxy for dose coverage, enabling effective learning without direct dose calculations. Spot-count representation establishes an avenue to employ deep learning to the task of EL pre-selection. (4) The unsupervised approach of this study not only eliminates the complications associated with ground truth generation, but also leaves objective function open-ended where multiple objectives can be designed and simultaneously optimized, paving the way to deep learning-based multi-criteria optimization of SPArc plan optimization. (5) SPArcdl enhances model inference speed by generating optimal EL sequences within 1 second, supporting fast clinical decision-making and planning workflows.
One limitation of SPArcdl is the instability of its predictions, which is largely attributed to the complexity involved in tuning the loss weight parameters. Future work will focus on developing an automated approach to replace the current manual and time-consuming process of objective weight tuning. Another limitation of this study stems from computational resource constraints on the research platform, which necessitated sampling the full-arc gantry angles at 5-degree intervals. We plan to enhance computational capacity to enable the analysis of a greater number of gantry angles in future investigations. Additionally, the current approach only incorporates spot counts for the clinical target volume (CTV) and brainstem, along with the mean water equivalent thickness (WET) value for each angle. Other potentially informative features—such as beamlet-level characteristics, geometric parameters, and a broader range of WET-related features [15]—were not considered. Incorporating these features as model inputs may lead to improved performance.
5 Conclusion
The proposed SPArcdl is able to achieve comparable or superior plan quality for delivery efficient proton arc therapy plan optimization, compared to the previous work SPArcps. The proposed spot-count representation establishes a foundation for unsupervised deep learning application in EL pre-selection of PAT plan optimization. We hope this work can inspire more deep learning-based methods in this field.